# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v2

@@ 6.3. VLM Quantization Results
In Figure 9, we compare 0-shot and multi-shot (4,8, 16 etc. in context examples) weight-only quantization performance of OpenFlamingo-9B on the COCO captioning task and VILA-7B on VizWiZ and TextVQA benchmarks. At W4A16 quantization, MicroScopiQ consistently outperforms all baselines and achieves on average less than 1% accuracy drop compared to the full-precision baseline, demonstrating the flexibility and widespread applicability of our method. Furthermore, at W2A16 quantization, MicroScopiQ still achieves high accuracy ($<4\%$ accuracy drop) and even outperforms several W4A16 quantization baselines.
We also quantitatively evaluate the performance of MicroScopiQ W2A16 quantized OpenFlamingo-9B VLM on the quality of generated captions for images drawn from the COCO captioning task in Figure 8. The MicroScopiQ quantized model generates accurate captions and preserves the overall word semantics compared to OliVe. For instance, in the second figure, OliVe mislabels the boat as a van, whereas MicroScopiQ is able to correctly identify it as a boat.
Table 5. Compute Area and Performance-Per-Unit-Area comparison.
  Architecture Component $\langle$ Area ($\mu m^{2}$), # Units $\rangle$ Total area ($\mu m^{2}$) Outlier support overhead Compute density (TOPS/$mm^{2}$)
@@
Figure 10. Iso-accuracy comparison between different accelerators. MicroScopiQ-v1 and v2 are executed on the same accelerator and correspond to two different data-precision distributions.
6.4. Accelerator Results
Area. In Table 5, we compare the accelerator area breakdown of MicroScopiQ with baselines GOBO and OliVe. For a fair evaluation, all accelerators have identical configurations with same number of PEs. MicroScopiQ has a very low compute overhead $19.76\%$, compared to OliVe’s $32.47\%$. This is because, units like ReCoN which perform three simple operations for outlier processing has a very low area utilization, 20$\sim$ 20$\times$ 20 × lower than having outlier-specific PEs like in GOBO. ReCoN is also time-multiplexed across all rows resulting in minimal area overhead (see section 6.5 for effects of time-multiplexing on performance). While OliVe has lower total area than GOBO, it has a higher compute-overhead because of complex encoding and decoding units. The simple INT operations of MicroScopiQ allows the packing of more compute power with minimal overheads. A similar multi-precision design in OliVe would require 3.5-4$\times$ higher PE area due to complex exponent-integer processing.
Maximum Performance Per Unit Area ($TOPS/mm^{2}$). We also estimate the maximum possible compute throughput per unit area of each accelerator, using LLaMA3-8B as the workload and report it as the compute density in Table 5. To achieve maximum throughput for MicroScopiQ, FMs must be quantized to low precision (e.g.. $bb=2$). MicroScopiQ achieves nearly 2×2$\times$2 × and 20×20$\times$20 × higher compute density than OliVe and GOBO respectively. This is because of higher total area of OliVe and single-precision MAC operation.
[…]
Iso-Accuracy Energy Comparison. In Figure 10(c), we show the normalized energy consumption of different accelerators, composed of static and dynamic energy. MicroScopiQ v2 has the lowest energy consumption. Compared to designs like GOBO, OLAccel and AdaptivFloat, where computations happen at higher-precisions in 8-bit and 32-bit PEs, MicroScopiQ v2 has a significant advantage in terms of the core and DRAM dynamic energy consumption with simple 2-/4-bit INT PEs. On average MicroScopiQ v2, has 1.5×1.5$\times$1.5 × lower energy consumption compared to all baselines across different FMs.
@@ 6.4. Accelerator Results
$+$ Quantize all weights to INT-4128 9.53 ($\downarrow 0.74$)
$+$ Quantize all weights to INT-2128 39.48 ($\uparrow 29.95$)
@@
$+$ Quantize activations to INT-8128, $\alpha=0.5$ 9.08 ($\uparrow 0.11$)
@@
6.5. Ablations and Discussions
Per-component accuracy impact. In Table 6 we examine the accuracy gained or lost by progressively incorporating the quantization techniques proposed in MicroScopiQ. We initially adopt INT-4 scalar quantization on the LLaMA3-8B full-precision baseline. We then apply group quantization with INT-4128. However, on decreasing precision to 2-bit INTs i.e., MX-INT-2, we observe a spike in PPL, due to higher outlier quantization error. Upon quantizing outliers to MX-FP-48,8, we regain the lost performance due to preserving outlier values at higher-precision and FP-format. Furthermore, we observe that pruning of least important weights causes a minor increase in PPL, which is regained by Hessian update.
@@ 6.5. Ablations and Discussions
Activation and KV-cache quantization. To demonstrate that the quantization techniques proposed in MicroScopiQ are complementary to activation quantization and KV-cache quantization, we show in Table 6 the effects of including these techniques. We follow simple group quantization for activations and handle activation outliers by migrating activation quantization difficulty to weights with a SmoothQuant-like migration strength $\alpha=0.5$ and 2-bit KV-cache quantization technique proposed in (Liu et al., 2024b). The K is quantized per channel, and the V is quantized per token with a MaB of size 128 and residual token length (R) (Liu et al., 2024b) of 128 for both key and value.
Enhanced low-bit quantization. Our method is orthogonal to the techniques proposed in OmniQuant. We can combine our method with OmniQuant to further improve the quantization performance of MicroScopiQ, advancing the SoTA and making it a more practical setting. Specifically, we employ OmniQuant’s Learnable Weight Clipping (LWC) to learn both the inlier and outlier scale factors (clipping threshold) of MicroScopiQ and leverage the Learnable Equivalent Transformation (LET) to migrate the quantization difficulty of activations to weights. In Table 4 we compare the performance of MicroScopiQ enhanced with OmniQuant against baseline OmniQuant. We employ the INT-(4/8)128 format for activation quantization. MicroScopiQ w/ OmniQuant outshines baselines across all settings, delivering improvements up to 22%.
Outlier group size. In Figure 11(a), we analyze the effect of different $B_{\mu}$ values on outlier diversity (standard deviation: red line), PPL (green bar), and EBW (blue line) for LLaMa3-8B. For $B_{\mu}=2,4$, accuracy degradation occurs due to outlier pruning, as many mBs have $\geq B_{\mu}/2$ outliers. As $B_{\mu}$ increases, outliers further apart show greater diversity, causing higher standard deviation. For $B_{\mu}\geq 32$, this results in increased quantization error and higher PPL, as MX-FP scales are shared across more diverse outlier values. Larger group sizes also significantly increase EBW due to metadata overhead. A balance is achieved at $B_{\mu}=8$.
@@
Analysis of time-multiplexed ReCoN. Firstly, inFigure 12(a), we show the layout of different variants of MicroScopiQ with varying number of ReCoN units. With the addition of more ReCoN units, fewer rows of PEs share a single ReCoN. When the number of ReCoN units equals the number of rows, each PE row gets a dedicated ReCoN unit. For a $64\times 64$ MicroScopiQ accelerator, we study the number of access conflicts for different number of ReCoN units in Figure 11(b). Evidently, the number of access conflicts to ReCoN which is measured as the percent of total number of accesses to ReCoN that result in conflicts is under 3% for all models. With the addition of more ReCoN units, the access conflicts tend to 0%. When accelerator area is a critical factor for an application, a single ReCoN unit is sufficient to reach good inference performance (See Figure 10). However, when latency is critical for an application, multiple ReCoN units can be employed with layout similar to as shown in Figure 12(a). Figure 14(a) demonstrates the impact of multiple ReCoN units on the compute area utilization and average latency for a LLaMA3-8B FM. Using upto 4 ReCoN units results in 18% improved latency with only 1.31×1.31$\times$1.31 × higher area. For practical systolic array sizes, 4/8 ReCoN units are sufficient to achieve ideal conflict-free performance.
6.6. Scalability and Timing Analysis
@@
Scalability. To study how MicroScopiQ scales to different PE array sizes, we compare the area against OliVe at three different scales: $8\times 8,16\times 16,128\times 128$ in Figure 13. As discussed above, since varying the number of ReCoN units trades off area for performance, we also compare three different versions of MicroScopiQ with 1,2 or 4 ReCoN units. For the single ReCoN variant of MicroScopiQ, we observe that as the design scales to higher array sizes, the area overhead of ReCoN significantly decreases, with the on-chip area being dominated by the large PE array and on-chip SRAM (weight, activation buffers etc.). We observe similar trends for 2 or 4 ReCoN units as well. Particularly at large PE array sizes such as $128\times 128$, adding 4 ReCoN units results in a minimal overhead of just 13%. This is because of two factors: (a) domination of on-chip area by PE array and (b) the column-wise bus from PEs to ReCoN can be routed over logic as the PE array scales. Therefore, at larger scales, it is beneficial to incorporate additional ReCoN units to perform high-throughput computation without conflicts to access ReCoN. OliVe at lower scales has a significantly higher area compared to MicroScopiQ and at higher scales has area comparable to MicroScopiQ. It is important to note that at all scales, MicroScopiQ has 23×2-3$\times$2 - 3 × higher throughput than OliVe (see Table 5).
Timing Analysis. MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. The timing critical path for MicroScopiQ is from the local weights registers to multipliers within the PE. The global column-wise bus (the longest wire in MicroScopiQ) from PEs to ReCoN is not on the critical path and therefore does not limit scalability  [^8] For extremely large hypothetical arrays, if global wire becomes a bottleneck, it can be resolved by pipelining/adding more ReCoN units (for timing) or laying wires over logic (for area)..
@@
Implementation overhead on NoC-based accelerators. The global column-wise bus is a design commonly seen in AI ASICs like (Chen et al., 2019; Firoozshahian et al., 2023; Shao et al., 2019; Tong et al., 2024). These global buses are used in conjunction with the NoC in these designs for data distribution, layout reordering. We modify two popular large-scale accelerators MTIA and Eyeriss-v2 that employ NoCs in their design to support a dedicated ReCoN and modify their PEs to fully integrate MicroScopiQ. As shown in Figure 14(b), compared to the baseline design of MTIA and Eyeriss-v2, integrating ReCoN and MicroScopiQ PEs results in only a $3\%$ and $2.3\%$ increase in compute area, respectively. Since these designs already have long wires and NoCs, integration overhead of MicroScopiQ is marginal. This demonstrates that the benefits achieved by integrating ReCoN are attained at minimal to no additional cost in real-world accelerators.
@@
Figure 14. (a) Effect of multiple time-multiplexed ReCoN units on the compute area and inference latency, (b) Implementation overhead of MicroScopiQ in existing NoC-based accelerators (Firoozshahian et al., 2023; Chen et al., 2019)
@@ 7. Related Work
Model compression for LLMs. LLM weight Compression scheme like quantization often rely on mixed-precision quantization with different outlier and inlier bit-widths(Lee et al., 2024; Dettmers et al., 2023; Zhao et al., 2024),(Huang et al., 2024). Other class of quantization rely on channel-wise shared scaling with different scale factors for outliers (Lee et al., 2024; Shao et al., 2023),(Huang et al., 2024). An orthogonal line of work on LLM quantization explores techniques to mitigate outliers from weight and activation distributions (Liu et al., 2024c), (Ashkboos et al., 2024b) to make quantization easier. Furthermore, recent work such as (Sharify et al., 2024) also explores the combined effect of multiple PTQ techniques for LLMs.
LLM weight pruning methods (Frantar and Alistarh, 2023; Yin et al., 2023) rely on approximate solvers to identify the importance of weights based on magnitude of weights or weights activation combined (Sun et al., 2023). Recently researchers have presented hardware friendly ordered pruning to trim model size (Ashkboos et al., 2024a). However, they fail to achieve satisfactory performance even at low sparsity levels(Lu et al., 2024).
Unified pruning and quantization. A recent work SDQ (Jeong et al., 2024), also explored the idea of how pruning can be effectively leveraged to complement outlier-aware quantization. Particularly, SDQ proposes to decompose a vector of weights into two separate inlier and outlier vectors each approximated and compressed into different N:M structured sparse patterns. Each vector is quantized individually, with the outliers quantized at twice the precision of inliers. MicroScopiQ can be considered as a generalized version of SDQ that does not impose a fixed N:M pattern and ensures hardware efficiency and aligned memory by ensuring a consistent bit-budget and data type per element. In a future version of this work, we shall conduct a detailed comparison against SDQ.
Accelerator for quantized models. Accelerators like GOBO (Zadeh et al., 2020) and OLAccel (Park et al., 2018) use high-precision PEs for outliers and low-precision PEs for inliers. OliVe (Guo et al., 2023) incorporates encoding/decoding units in a systolic array for inlier-outlier formats, while other accelerators (Tambe et al., 2020; Ramachandran et al., 2024b; Guo et al., 2022) propose adaptive formats and hybrid FP PEs. However, they often face issues like unaligned memory access and large overheads.
NoCs in ML Accelerators. NoCs are widely used in accelerators: (Qin et al., 2020; Muñoz-Martínez et al., 2023) employ a Benes NoC for flexible data distribution, (Tong et al., 2024) uses a butterfly NoC for layout transformation, and (Firoozshahian et al., 2023; Chen et al., 2019) utilize mesh NoCs for data delivery. ReCoN leverages a butterfly NoC for efficient processing.

### v3

@@ 6.3. VLM Quantization Results
In Figure 8, we compare 0-shot and multi-shot (4,8, 16, 32) weight-only quantization performance of OpenFlamingo-9B on the COCO captioning task and VILA-7B on VizWiZ, TextVQA benchmarks. At W4A16 quantization, MicroScopiQ consistently outperforms all baselines and achieves on average less than 1% accuracy drop compared to the full-precision baseline, demonstrating the flexibility and widespread applicability of our method. Furthermore, at W2A16 quantization, MicroScopiQ achieves high accuracy ($<4\%$ accuracy drop), outperforming several W4A16 baselines.
6.4. Accelerator Results
Compute Area. In Table 5, we compare the accelerator compute area breakdown of MicroScopiQ (w/ 1 ReCoN unit) with baselines GOBO and OliVe for a $64\times 64$ array design. For a fair evaluation, all accelerators have identical configurations with same number of PEs. MicroScopiQ has a very low compute overhead $8.63\%$, compared to OliVe’s $9.90\%$. This is because, units like ReCoN which perform three simple operations for outlier processing has a very low area utilization compared to OliVe’s encoders and decoders. ReCoN is also time-multiplexed across all rows resulting in minimal area overhead (see section 6.5). Furthermore. the simple INT operations of MicroScopiQ allows the packing of more compute power with minimal overheads. A similar multi-precision design in OliVe would require 3.5-4$\times$ higher multi-precision support area due to complex exponent-integer PEs. GOBO has an order of magnitude higher compute area compared to OliVe and MicroScopiQ due to large per-PE area.
Table 5. Compute area and density comparison for a $64\times 64$ array at 7nm process technology.
  Architecture Component $\langle$ Area ($\mu m^{2}$), # Units $\rangle$ Compute area ($mm^{2}$) Compute overhead Compute density (TOPS/$mm^{2}$)
@@ 6.4. Accelerator Results
Figure 9. Iso-accuracy comparison of different accelerators. MicroScopiQ-v1 and -v2 are executed on the same accelerator and correspond to two different data-precision distributions.
Maximum Performance Per Unit Area ($TOPS/mm^{2}$). We also estimate the peak compute throughput per unit area of each accelerator, using LLaMA3-8B as the workload and report it as the compute density in Table 5. To achieve peak throughput for MicroScopiQ, FMs must be quantized to low precision (e.g.. $bb=2$). MicroScopiQ achieves nearly 2$\times$ and 14$\times$ higher compute density than OliVe and GOBO respectively.
@@
Iso-Accuracy Energy Comparison. In Figure 9(c), we show the normalized energy consumption of different accelerators, composed of static and dynamic energy. MicroScopiQ v2 has the lowest energy consumption. Compared to designs like GOBO, OLAccel and AdaptivFloat, where computations happen at higher-precisions in 8-bit and 32-bit PEs, MicroScopiQ v2 has a significant advantage in terms of the core and DRAM dynamic energy consumption with simple INT PEs. On average MicroScopiQ v2, has 1.5$\times$ lower energy consumption compared to all baselines across different FMs.
@@
$+$ Quantize all weights to MX-INT-4128 9.53 ($\downarrow 0.74$)
$+$ Quantize all weights to MX-INT-2128 39.48 ($\uparrow 29.95$)
@@
$+$ Quantize activations to MX-INT-8128, $\alpha=0.7$ 9.08 ($\uparrow 0.11$)
@@
6.5. Algorithm ablations
Per-component accuracy impact. In Table 6 we examine the accuracy gained or lost by progressively incorporating the quantization techniques proposed in MicroScopiQ. We initially adopt INT-4 scalar quantization on the LLaMA3-8B full-precision baseline. We then apply MX-INT-4128. However, on decreasing precision to 2-bit INTs i.e., MX-INT-2128, we observe a spike in PPL, due to higher outlier quantization error. Upon quantizing outliers to MX-FP-48,8, we regain the lost performance due to preserving outlier values at higher-precision and FP-format. Furthermore, we observe that pruning of least important weights causes a minor increase in PPL, which is regained by Hessian update. We observe that quantizing activations with simple MX-INT-8128 with a high $\alpha$ results in a very minimal increase in PPL due to the robustness of MicroScopiQ.
KV-cache quantization. In Table 6 we also report the impact of KV-cache quantization on performance. Following the 2-bit KV-cache quantization technique proposed in (Liu et al., 2024b) we quantize K per channel, and the V per token with a MaB of size 128 and residual token length (R) (Liu et al., 2024b) of 128 for both key and value.
Omni-MicroScopiQ. Our method is orthogonal to the techniques proposed in OmniQuant. We combine our method with OmniQuant to further improve the quantization performance of MicroScopiQ. We employ OmniQuant’s Learnable Weight Clipping (LWC) to learn both the inlier and outlier scale factors of MicroScopiQ and leverage the Learnable Equivalent Transformation (LET) to learn to migrate the quantization difficulty of activations to weights. In Table 4 we compare the performance of MicroScopiQ enhanced with OmniQuant (Omni-MicroScopiQ) against OmniQuant. Omni-MicroScopiQ delivers improvements up to 22%.
@@ 6.5. Algorithm ablations
Outlier group size. In Figure 10(a), we analyze the effect of different $B_{\mu}$ values on outlier diversity (measure by the standard deviation: red line), PPL (green bar), and EBW (blue line) for LLaMa3-8B. For $B_{\mu}=2,4$, accuracy degradation occurs due to outlier pruning, as many mBs have $\geq B_{\mu}/2$ outliers. As $B_{\mu}$ increases, outliers further apart show greater diversity, evidenced by the higher standard deviation. For $B_{\mu}\geq 32$, this results in increased quantization error and higher PPL, as MX-FP scales are shared across more diverse outlier values. Larger group sizes also significantly increase EBW due to metadata overhead. A balance is achieved at $B_{\mu}=8$.
@@
6.6. Analysis of time-multiplexed ReCoN
Accelerator design variants. InFigure 11(a), we show the layout of different variants of MicroScopiQ with varying number of ReCoN units, each of which can be employed based on application requirements. For area critical applications, design A, with a single ReCoN can be employed to have minimal area utilization with a slight increase in latency due to access conflicts to ReCoN. With the addition of more ReCoN units, fewer rows of PEs share a single ReCoN, dramatically reducing access conflicts to the shared ReCoN unit (design B). When the number of ReCoN units equals the number of rows, each PE row gets a dedicated ReCoN unit as shown in design C, which can be employed for latency critical applications. As we shall demonstrate subsequently, we identified that in up to a $128\times 128$ PE array size, a total of 8 ReCoN units are sufficient to achieve peak performance with zero access conflicts to ReCoN.
Optimal # of ReCoN units. For a $64\times 64$ MicroScopiQ accelerator, we study the number of access conflicts for different number of ReCoN units in Figure 10(b). Evidently, the number of access conflicts to ReCoN which is measured as the percent of total number of accesses to ReCoN that result in conflicts is under 3%. With the progressive addition of more ReCoN units, the access conflicts tend to 0%. In Figure 9, despite the access conflicts, 1 ReCoN unit is sufficient to reach good inference performance in an iso-accuracy scenario. For latency critical applications, in Figure 13(a), we quantify the performance improvement and impact on compute area by incorporating multiple ReCoN units for a LLaMA3-8B FM. Up to 8 ReCoN units enables peak performance with 21% improved latency and only 1.58×1.58$\times$1.58 × higher compute area. For large array sizes, up to 8 ReCoN units are sufficient for peak performance.
6.7. Scalability and Timing Analysis
@@
Scalability. To study how MicroScopiQ scales to different PE array sizes, we compare the total area (including on-chip weight and iAct/oAct buffers) against OliVe at three different scales: $8\times 8,16\times 16,128\times 128$ in Figure 12. Following (Sumbul et al., 2022), for $8\times 8$ array we employ 16 kB iAct and oAct buffers and 32 kB weight buffer. We progressively scale the buffers for larger arrays to ensure full array utilization. Since varying the number of ReCoN units trades off area for performance, we also compare three versions of MicroScopiQ with 1, 2, 8 ReCoN units. For the single ReCoN variant of MicroScopiQ, as the design scales to higher array sizes, the area overhead of ReCoN decreases ($128\times 128$ has overhead of 3%), with the on-chip area being dominated by the large PE array and buffers. We observe similar trends for 2 or 8 ReCoN units. At large PE array sizes such as $128\times 128$, adding 8 ReCoN units results in a minimal overhead of just 11%. Particularly at large array sizes the column-wise bus from PEs to ReCoN can be routed over logic as the PE array scales. With same buffer configurations, OliVe has a higher area compared to a single ReCoN MicroScopiQ variant, while comparable area to MicroScopiQ with 8 ReCoN units. At all variants, MicroScopiQ has 23×2-3$\times$2 - 3 × higher performance than OliVe (see Table 5).
@@
Timing Analysis. MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. The critical path for MicroScopiQ is from the local weights registers to multipliers within the PE. The column-wise bus (the longest wire in MicroScopiQ) from PEs to ReCoN is not on the critical path and therefore does not limit scalability  [^6] For extremely large arrays, if global wire becomes a bottleneck, it can be resolved by pipelining/adding more ReCoN units (for timing) or laying wires over logic (for area).. Implementation overhead on NoC-based accelerators. The global column-wise bus is a design commonly seen in AI ASICs like (Chen et al., 2019; Firoozshahian et al., 2023; Shao et al., 2019; Tong et al., 2024). These global buses are used in conjunction with the NoC in these designs for data distribution, layout reordering (Tong et al., 2024). We modify two large-scale accelerators MTIA and Eyeriss-v2 that employ NoCs in their design to support MicroScopiQ, by incorporating ReCoN functionality in the accelerator NoCs and modifying PEs to support all MicroScopiQ specific operations. As shown in Figure 13(b), compared to the baseline design of MTIA and Eyeriss-v2, integrating ReCoN and MicroScopiQ PEs results in only a $3\%$ and $2.3\%$ increase in compute area, respectively. Since these designs already employ NoCs, integration overhead of MicroScopiQ is minimal. This demonstrates that the benefits achieved by MicroScopiQ are attained at minimal to no additional cost.
@@
Figure 13. (a) Effect of time-multiplexed ReCoN units on the compute area and inference latency, (b) Implementation overhead of MicroScopiQ in NoC-based accelerators (Firoozshahian et al., 2023; Chen et al., 2019)
@@ 7. Related Work
Model compression for LLMs. LLM weight Compression scheme like quantization often rely on mixed-precision quantization with different outlier and inlier bit-widths(Lee et al., 2024; Dettmers et al., 2023; Zhao et al., 2024; Jeong et al., 2024; Huang et al., 2024). For instance, SDQ (Jeong et al., 2024) decomposes a vector of weights into two separate inlier and outlier vectors each quantized at different precisions. Other class of quantization rely on channel-wise shared scaling with different scale factors for outliers (Lee et al., 2024; Shao et al., 2023; Huang et al., 2024). Other quantization techniques mitigate outliers from weights and activations (Liu et al., 2024c; Xiao et al., 2023; Ashkboos et al., 2024b) to make quantization easier.
LLM weight pruning methods (Frantar and Alistarh, 2023; Yin et al., 2023; Ashkboos et al., 2024a) rely on approximate solvers to identify the importance of weights based on magnitude of weights or weights activation combined (Sun et al., 2023). However, they fail to achieve satisfactory performance even at low sparsity levels(Lu et al., 2024). MicroScopiQ provides a unified solution for outlier-aware quantization by enabling flexible pruning of least-important weights to distribute additional outlier bits, thereby, simultaneously achieving high accuracy and hardware efficiency.
Accelerator for quantized models. Accelerators like GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) use high-precision PEs for outliers and low-precision PEs for inliers. OliVe (Guo et al., 2023) incorporates encoding/decoding units in a PE array for inlier-outlier formats, while other accelerators (Tambe et al., 2020; Ramachandran et al., 2024b; Guo et al., 2022) propose adaptive formats and hybrid FP PEs. However, they often suffer from unaligned memory access and large overheads. MicroScopiQ accelerator introduces a novel low overhead NoC called ReCoN, that effectively abstracts FP-format complexity from INT-PEs
NoCs in ML Accelerators. NoCs are widely used in accelerators: (Qin et al., 2020; Muñoz-Martínez et al., 2023) employ a Benes NoC for data distribution, (Tong et al., 2024) uses a butterfly NoC for layout transformation, and (Firoozshahian et al., 2023; Chen et al., 2019) utilize mesh NoCs. ReCoN is a butterfly NoC employed to compute FP partial sums.